Using the Formula n=1/(ME)2 to Determine Sample Size

When we want 95 % confidence and use the conservative estimate of p = 0.5 , we can use the simple formula n = 1 M ⁢ E 2 to roughly determine the sample size needed for a given margin of error M ⁢ E . Use this formula to determine the sample size needed for a margin of error of 0.04 .

Enter the exact answer.
Answer #1

we have given margin of error = 0.04

n=1/(ME)2

n = 1/(0.04)2

n =1/0.0016

n= 625

Sample Size = 625